OpenStudy (clalgee):

An original data set has the values 14, 2, 11, 6, and 2. What would need to be added to each value to change the mean to 15? a. 1 b. 8 c. 9 d. 13 Please help! Only one question :-/

OpenStudy (amistre64):

what is the mean of the current data set?

OpenStudy (amistre64):

\[mean=\frac{14+2+11+6+2}{5}\] \[mean+k=\frac{(14+k)+(2+k)+(11+k)+(6+k)+(2+k)}{5}\] \[mean+k=\frac{14+2+11+6+2+5k}{5}\] \[mean+k=\frac{14+2+11+6+2}{5}+\frac{5k}{5}\] \[mean+k=\frac{14+2+11+6+2}{5}+k\]

OpenStudy (amistre64):

in other words, the mean of similar (dare i say congruent) data sets differ by a constant

OpenStudy (amistre64):